# Dual Frequency Branch Framework with Reconstructed Sliding Windows Attention for AI-Generated Image Detection

### v2

@@ IV-A Dataset
AIGCDetectBenchmark [48]: This dataset comprehensively includes GAN and diffusion, and introduces some new diffusion generation methods as detection datasets, including ProGAN, StyleGAN, StyleGAN2, BigGAN, CycleGAN, StarGAN, GauGAN, Deepfake, WFIR, Midjourney, SDv1.4, SDv1.5, ADM, GLIDE, Wukong, VQDM, which is a huge challenge for deepfake detection.
AIGIBench [49]: AIGIBench is the first benchmark to conduct a comprehensive evaluation of four critical components in the AIGI detection pipeline, and comprehensively simulates state-of-the-art image generation methods, including: (a) GAN-based noise-to-image generation (ProGAN, StyleGAN3, StyleGAN-XL, StyleSwim, R3GAN, and WFIR), (b) Diffusion for text-to-image generation (SD-XL, SD-3, DALLE-3, Midjourney-v6, FLUX.1-dev, Imagen-3, and GLIDE), (c) GANs for deepfake (BlendFace, E4S, FaceSwap, InSwap, and SimSwap), and (d) Diffusion for personalized generation (InstantID [27], Infinite-ID [28], PhotoMaker [29], BLIP-Diffusion [30], and IP-Adapter [31]). It also includes 2 general subsets featuring fake images collected from social media platforms, which called CommunityAI and SocialRF.
@@ IV-C Quantitative analysis
We compare with the previous methods: CNNDetection [50], Patchfor [34], F3Net [37], FrePGAN [7], GANDetection [57], LGrad [8], CLIPDetection [43], FreqNet [9] ,NPR [3], FatFormer [15], Ladeda [35], C2P_CLIP [17], AIDE [18]. We conduct comprehensive experiments on five datasets to verify the effectiveness of our method. Specifically, the results of Patchfor [34], F3Net [37], FrePGAN [7], GANDetection [57] all come from NPR [3], and we obtain results of FatFormer [15], C2P_CLIP [17] using the official pre-trained model. In addition, we retrain all other methods following the paradigms of [43], which adopts specific 4-class training settings, denoted as (car, cat, chair, horse).
@@
GANGen-Detection[9]. Since our method uses ProGAN for training, we first perform cross-domain testing on GAN. As shown in Table I, our detection on GAN has achieved good results. Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 97.3% and 98.8%, respectively. Compared to Ladeda and C2P_CLIP, our method improves mean Acc. by 0.1% and 0.9%, respectively. However, on Fatformer, the accuracy is 1.1% higher than our method, which needs to be improved.
@@
DiffusionForensics [46] & UniversalFakeDetect [43]. To further evaluate the generalization ability of our method, we train on ProGAN and test on images generated by various diffusion models. The results are presented in Tables II and III. Despite being trained exclusively on ProGAN-generated images, our method demonstrates strong generalization capabilities across diverse diffusion models. Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 98.8% and 100.0%, respectively. Compared with the best detection methods NPR and Ladeda, our method improves mean Acc. by 3.0% and 3.7%, respectively. However, methods Fatformer, C2P_CLIP, and AIDE, which have good results on other datasets, have an accuracy rate of less than 90% on this dataset, and even C2P_CLIP has only 80.7% accuracy. The most direct speculation for this result is that they used CLIP in the network skeleton, but the real reason is still worth further exploration. Additionally, on the UniversalFakeDetect dataset [43], our method achieves an average accuracy of 96.4%, 2.2% accuracy improvement over the best method.
@@
GenImage [48]. Compared with testing only on diffusion or GAN, GenImage not only includes the classic GAN generation method, but also adds new and more realistic diffusion generation datasets, such as DALLE2. The test results are shown in Table IV. Our method achieves a notable improvement in mean accuracy, outperforming C2P_CLIP and Ladeda by 2.4% and 4.0%, respectively, and reaching an accuracy of 90.6%.
@@
AIGIBench [49]. AIGIBench proposes a more challenging forgery detection dataset to evaluate the performance of forgery detection models from different perspectives. The test results are shown in Table V, our detection has achieved good results. Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 77.9% and 82.0%, respectively. Compared to Ladeda and C2P_CLIP, our method improves mean Acc. by 3.0% and 5.9%, respectively. Although our results are the best, there are still major problems and improvements: 1) Although our method has achieved an accuracy rate of more than 90% in most GAN and diffusion scenarios, it still has an accuracy rate of only about 60% in the dataset collected in unknown generated social scenarios. Other methods also have very low accuracy rates, which still needs further research and improvement. 2) When it comes to face-swapping forgery detection, the performance is generally low, even less than 50%. Almost all methods have this problem, which is also a problem that needs to be considered and improved.
@@
IV-D Qualitative Analysis
IV-E Module Ablation Studies.
FFT alone has negative impacts on Attention. When we remove our main module (DWT-Based Window Tiling), the performance of our method drops significantly, falling below 50%. In our design, FFT is integrated as a parallel branch to complement DWT, leveraging its phase information to capture long-range dependencies that enhance fine-grained feature richness for reconstructed sliding window attention. However, the localized focus of RSWAttention, optimized for fine-grained details, appears incompatible with FFT’s global frequency representation when not moderated by DWT’s multi-scale decomposition or Window Tiling’s spatial structuring. This misalignment suggests that FFT alone introduces overly coarse features, diluting the attention mechanism’s ability to discern subtle, localized artifacts critical for deepfake identification.
It is effective to use the phase part of FFT as another branch.As shown in Table VI, when the phase part of FFT is excluded, the average accuracy of the model decreases 1.8%, indicating that FFT enhances the attention mechanism’s capacity to model interdependencies between distant image regions. FFT, as another frequency domain branch, enriches the local features within the window attention from different perspectives.
IV-F Comparative Studies of Module Design
@@
1) The reconstructed sliding window attention mechanism greatly improves the ability of attention to capture forgery features. As shown in Table VII, compared with WAttention, our RSWAttention improves the mean Acc. and mean A.P. by 24.2% and 13.1% respectively, which suggests that the RSWAttention greatly promotes the attention mechanism to capture fine-grained forgery traces within the local window while modeling the importance and interdependence between local features.
@@
2) Utilizing only the phase component of the FFT can reduce features not related to forgery compared to using the full FFT. As shown in Table VIII, if we use not only the phase part of the FFT but also the amplitude part as input to the network, the average accuracy of the model drops by 5.2%, confirming that amplitude introduces noise or irrelevant intensity information that confounds the detector. In contrast, the phase, encoding structural details like edges and shapes, aligns with the artifact-relevant features needed for detection. This finding corroborates prior image processing research, where phase information is prioritized over amplitude for structural analysis, reinforcing our design choice to leverage FFT’s phase as a complementary input branch.
@@
3) All subbands of DWT are indispensable. Each frequency domain subband of DWT has its own unique meaning, the low-frequency part retains the overall structure and semantic background of the image, and the high-frequency part captures subtle texture and edge information. Deleting any subband will reduce the comprehensiveness of the information in the window. As shown in Table IX, whether using only one frequency domain subband or omitting some frequency domain subbands, the performance of the model will be degraded.
IV-G Hyperparameter Ablation Study.

### v3

@@ IV-A Dataset
AIGCDetectBenchmark [83]: This dataset comprehensively includes GAN and diffusion, and introduces some new diffusion generation methods as detection datasets, including ProGAN, StyleGAN, StyleGAN2, BigGAN, CycleGAN, StarGAN, GauGAN, Deepfake, WFIR, Midjourney, SDv1.4, SDv1.5, ADM, GLIDE, Wukong, VQDM, which is a huge challenge.
AIGIBench [35]: AIGIBench is the first benchmark to conduct a comprehensive evaluation of four critical components in the AIGI detection pipeline, and comprehensively simulates state-of-the-art image generation methods, including: (a) GAN-based noise-to-image generation (ProGAN, StyleGAN3, StyleGAN-XL, StyleSwim, R3GAN, and WFIR), (b) Diffusion for text-to-image generation (SD-XL, SD-3, DALLE-3, Midjourney-v6, FLUX.1-dev, Imagen-3, and GLIDE), (c) GANs for deepfake (BlendFace, E4S, FaceSwap, InSwap, and SimSwap), and (d) Diffusion for personalized generation (InstantID [65], Infinite-ID [70], PhotoMaker [31], BLIP-Diffusion [29], and IP-Adapter [78]). It also includes 2 general subsets collected from social media platforms, which called CommunityAI and SocialRF.
@@ IV-B Implementation Details
| CNN-Spot | 54.9 | 66.6 | 54.2 | 76.0 | 53.3 | 72.9 | 53.0 | 71.3 | 51.9 | 63.7 | 52.0 | 64.5 | 51.6 | 63.1 | 51.8 | 61.3 | 52.8 | 67.4 |
| UnivFD | 69.7 | 87.6 | 79.1 | 99.8 | 77.9 | 94.5 | 78.5 | 95.3 | 95.0 | 99.3 | 99.0 | 99.3 | 74.0 | 92.5 | 87.3 | 97.5 | 82.6 | 95.7 |
| C2P-CLIP | 69.1 | 94.1 | 95.3 | 99.8 | 96.1 | 99.8 | 95.3 | 99.8 | 99.3 | 100.0 | 99.3 | 100.0 | 97.3 | 99.8 | 98.6 | 99.1 | 93.8 | 99.2 |
| CNN-Spot | 97.9 | 99.9 | 66.1 | 93.5 | 67.0 | 97.1 | 55.9 | 80.5 | 70.1 | 97.0 | 76.5 | 91.5 | 72.9 | 93.2 | 51.7 | 66.5 | 62.1 | 89.4 |
| UnivFD | 100.0 | 100.0 | 99.7 | 100.0 | 75.7 | 97.8 | 95.1 | 99.3 | 84.4 | 97.1 | 98.7 | 99.8 | 99.9 | 99.4 | 67.4 | 82.0 | 64.2 | 68.7 |
| C2P-CLIP | 100.0 | 100.0 | 99.2 | 100.0 | 95.6 | 99.9 | 99.1 | 100.0 | 96.4 | 99.5 | 97.3 | 100.0 | 99.6 | 100.0 | 93.8 | 98.6 | 94.8 | 99.5 |
| CNN-Spot | 52.1 | 67.8 | 51.7 | 67.1 | 51.7 | 64.7 | 50.6 | 59.2 | 50.6 | 59.2 | 51.5 | 63.5 | 50.5 | 58.5 | 50.1 | 49.2 | 61.3 | 76.7 |
| UnivFD | 67.6 | 87.2 | 61.7 | 85.5 | 56.7 | 74.5 | 63.1 | 87.2 | 62.8 | 86.3 | 84.9 | 96.8 | 71.2 | 90.9 | 49.5 | 60.2 | 73.5 | 83.6 |
| C2P-CLIP | 77.2 | 95.8 | 88.5 | 98.8 | 59.2 | 83.0 | 82.8 | 97.3 | 82.8 | 97.2 | 87.2 | 98.4 | 80.1 | 95.5 | 65.1 | 92.9 | 88.2 | 92.6 |
@@ IV-C Quantitative analysis
We compare with the previous methods: CNN-Spot [66], Patchfor [5], F3Net [47], FrePGAN [23], GANDetection [42], LGrad [57], UnivFD [45], FreqNet [55] ,NPR [56], FatFormer [37], Ladeda [4], C2P-CLIP [54], AIDE [74]. We conduct comprehensive experiments on five datasets to verify the effectiveness of our method. Specifically, the results of Patchfor [5], F3Net [47], FrePGAN [23], GANDetection [42] all come from NPR [56], and we obtain results of FatFormer [37], C2P-CLIP [54] using the official pre-trained model. In addition, we retrain all other methods following the paradigms of [45], which adopts specific 4-class training settings, denoted as (car, cat, chair, horse).
@@
GANGen-Detection[55]. Since our method uses ProGAN for training, we first perform cross-domain testing on GAN. As shown in Table I, our detection on GAN has achieved good results. Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 97.3% and 98.8%, respectively. Compared to Ladeda and C2P-CLIP, our method improves mean Acc. by 0.1% and 0.9%, respectively. However, on Fatformer, the accuracy is 1.1% higher than our method, which needs to be improved.
@@
DiffusionForensics [67] & UniversalFakeDetect [45]. To further evaluate the generalization ability of our method, we train on ProGAN and test on images generated by various diffusion models. The results are presented in Tables II and III. Despite being trained exclusively on ProGAN-generated images, our method demonstrates strong generalization capabilities across diverse diffusion models. Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 98.8% and 100.0%, respectively. Compared with the best detection methods NPR and Ladeda, our method improves mean Acc. by 3.0% and 3.7%, respectively. However, methods Fatformer, C2P-CLIP, and AIDE, which have good results on other datasets, have an accuracy rate of less than 90% on this dataset, and even C2P-CLIP has only 80.7% accuracy. The most direct speculation for this result is that they used CLIP in the network skeleton, but the real reason is still worth further exploration. Additionally, on the UniversalFakeDetect dataset [45], our method achieves an average accuracy of 96.4%, 2.2% accuracy improvement over the best method.
@@
GenImage [83]. Compared with testing only on diffusion or GAN, GenImage not only includes the classic GAN generation method, but also adds new and more realistic diffusion generation datasets, such as DALLE2. The test results are shown in Table IV. Our method achieves a notable improvement in mean accuracy, outperforming C2P-CLIP and Ladeda by 2.4% and 4.0%, respectively, and reaching an accuracy of 90.6%.
@@
AIGIBench [35]. AIGIBench proposes a more challenging forgery detection dataset to evaluate the performance of forgery detection models from different perspectives. The test results are shown in Table V, our detection has achieved good results. Specifically, it achieves mean Accuracy (Acc.) and mean Average Precision (A.P.) of 77.9% and 82.0%, respectively. Compared to Ladeda and C2P-CLIP, our method improves mean Acc. by 3.0% and 5.9%, respectively. Although our results are the best, there are still major problems and improvements: 1) Although our method has achieved an accuracy rate of more than 90% in most GAN and diffusion scenarios, it still has an accuracy rate of only about 60% in the dataset collected in unknown generated social scenarios. Other methods also have very low accuracy rates, which still needs further research and improvement. 2) When it comes to face-swapping forgery detection, the performance is generally low, even less than 50%. Almost all methods have this problem, which is also a problem that needs to be improved.
| CNN-Spot | 97.8 | 100.0 | 49.9 | 48.4 | 64.0 | 91.6 | 49.7 | 60.0 | 58.0 | 88.9 | 62.1 | 89.4 | 49.3 | 34.2 | 49.6 | 35.2 | 50.0 | 44.3 |
| UnivFD | 99.5 | 100.0 | 93.0 | 98.6 | 68.4 | 79.0 | 87.1 | 95.0 | 94.5 | 99.3 | 91.0 | 97.8 | 47.2 | 42.7 | 72.4 | 82.5 | 69.3 | 79.4 |
| C2P-CLIP | 100.0 | 100.0 | 95.4 | 99.8 | 93.8 | 98.5 | 94.1 | 98.1 | 95.2 | 100.0 | 94.8 | 99.5 | 55.2 | 60.5 | 69.8 | 76.1 | 73.5 | 82.2 |
| CNN-Spot | 50.1 | 44.3 | 49.7 | 45.3 | 50.3 | 50.3 | 49.9 | 39.3 | 51.5 | 67.3 | 50.4 | 53.8 | 51.0 | 50.3 | 52.8 | 66.3 | 52.6 | 65.5 |
| UnivFD | 48.9 | 48.2 | 50.1 | 51.5 | 46.3 | 39.6 | 44.8 | 35.6 | 69.4 | 81.7 | 46.8 | 38.4 | 48.1 | 44.1 | 49.0 | 48.6 | 50.7 | 53.6 |
| C2P-CLIP | 59.5 | 69.0 | 67.9 | 78.5 | 47.6 | 46.8 | 51.2 | 50.2 | 85.1 | 93.4 | 64.4 | 75.4 | 45.2 | 46.5 | 53.5 | 65.6 | 59.3 | 76.0 |
| CNN-Spot | 49.8 | 44.7 | 50.8 | 55.8 | 50.1 | 67.1 | 50.0 | 50.1 | 49.6 | 51.5 | 50.3 | 46.3 | 49.9 | 50.5 | 53.6 | 57.6 |
| UnivFD | 57.8 | 68.9 | 72.4 | 84.2 | 67.4 | 79.5 | 62.3 | 73.9 | 47.0 | 41.4 | 49.2 | 46.2 | 49.9 | 50.0 | 63.3 | 66.4 |
| C2P-CLIP | 84.9 | 93.8 | 85.2 | 93.4 | 86.8 | 93.9 | 55.1 | 65.9 | 50.6 | 52.0 | 56.4 | 62.3 | 51.0 | 54.5 | 71.0 | 77.3 |
IV-D Qualitative Analysis
IV-E Feature Visualization of Dual Frequency Branches
To further validate that the two frequency branches capture complementary artifact information, we visualize the features extracted by each branch, as shown in Figure 5. Specifically, for the DWT-based window tiling branch, the reconstructed images exhibit intensified edges and subtle darkening around object contours, indicating that this branch focuses on fine-grained local artifacts and high-frequency inconsistencies. In contrast, the FFT-based phase complement branch yields patch-like residual patterns that spread over large regions rather than aligning with specific edges, which arise from phase perturbations in the frequency domain and reflect global structural irregularities that are characteristic of different generators. Together, these observations support the effectiveness of our method in extracting artifact cues from complementary perspectives.
@@ IV-E Feature Visualization of Dual Frequency Branches
| Method | JPEG Compression | Gaussian Blur | ||||
| QF=95 | QF=85 | QF=75 | ||||
| CNN-Spot | 59.0 | 59.9 | 60.5 | 62.4 | 63.2 | 63.3 |
| UnivFD | 76.4 | 74.2 | 73.8 | 80.2 | 76.8 | 76.3 |
| NPR | 72.8 | 70.7 | 69.2 | 80.6 | 79.6 | 81.0 |
| LaDeDa | 74.4 | 70.4 | 68.0 | 84.8 | 83.2 | 83.5 |
| AIDE | 74.6 | 74.1 | 70.3 | 88.6 | 75.8 | 81.1 |
| Ours | 78.8 | 76.6 | 75.2 | 92.0 | 88.4 | 87.9 |
| Method | JPEG Compression | Gaussian Blur | ||||
| QF=95 | QF=85 | QF=75 | ||||
| CNN-Spot | 53.8 | 54.3 | 54.5 | 56.5 | 56.5 | 56.6 |
| UnivFD | 65.6 | 64.3 | 62.7 | 72.2 | 68.6 | 68.6 |
| NPR | 59.0 | 58.9 | 58.6 | 62.9 | 62.0 | 62.9 |
| LaDeDa | 61.6 | 60.0 | 58.2 | 64.4 | 66.0 | 65.6 |
| AIDE | 61.0 | 59.6 | 55.9 | 75.5 | 63.4 | 65.2 |
| Ours | 68.9 | 67.2 | 65.2 | 75.5 | 74.6 | 74.2 |
@@
IV-F Module Ablation Studies
FFT alone has negative impacts on Attention. When we remove our main module (DWT-based Window Tiling), the performance of our method drops significantly, falling below 50%. In our design, FFT is integrated as a parallel branch to complement DWT, leveraging its phase information to capture long-range dependencies that enhance fine-grained feature richness for reconstructed sliding window attention. However, the localized focus of RSWAttention, optimized for fine-grained details, appears incompatible with FFT’s global frequency representation when not moderated by DWT’s multi-scale decomposition or Window Tiling’s spatial structuring. This misalignment suggests that FFT alone introduces overly coarse features, diluting the attention mechanism’s ability to discern subtle, localized artifacts critical for detection.
It is effective to use the phase part of FFT as another branch. As shown in Table VI, when the phase part of FFT is excluded, the accuracy of model decreases 1.8%, indicating that FFT enhances the attention mechanism’s capacity to model interdependencies between distant image regions. FFT, as another frequency domain branch, enriches the local features within the window attention from different perspectives.
| Complement Branch | GANGen-Detection | DiffusionForensics | UniversalFakeDetect | GenImage | AIGIBench | mean Acc. / A.P. | |
| DCT | 95.9 / 98.1 | 97.2 / 100.0 | 95.8 / 99.0 | 87.9 / 93.3 | 75.7 / 80.1 | 90.5 / 92.5 | |
| FFT | 96.6 / 98.8 | 98.8 / 100.0 | 96.4 / 99.3 | 90.6 / 93.6 | 77.9 / 82.0 | 92.1 / 94.1 |
@@ IV-F Module Ablation Studies
| Window Tiling | DWT-Based | FFT-Based | GANGen-Detection | DiffusionForensics | UniversalFakeDetect | GenImage | AIGIBench | mean Acc. / A.P. |
| 92.5 / 98.2 | 83.8 / 99.8 | 93.8 / 99.0 | 87.4 / 93.1 | 71.4 / 77.1 | 85.4 / 92.3 | |||
| 95.8 / 98.8 | 91.1 / 99.9 | 95.9 / 99.3 | 89.2 / 92.2 | 73.8 / 76.7 | 88.3 / 92.0 | |||
| 92.9 / 97.0 | 84.0 / 99.2 | 89.4 / 97.7 | 82.1 / 91.0 | 64.6 / 72.1 | 81.1 / 89.5 | |||
| 74.4 / 80.7 | 70.4 / 89.7 | 66.4 / 73.1 | 67.4 / 72.6 | 59.6 / 66.2 | 68.6 / 76.4 | |||
| 95.5 / 98.3 | 91.9 / 99.9 | 94.8 / 99.0 | 88.9 / 91.9 | 71.8 / 74.7 | 87.4 / 91.3 | |||
| 96.6 / 98.8 | 98.8 / 100.0 | 96.4 / 99.3 | 90.6 / 93.6 | 77.9 / 82.0 | 92.1 / 94.1 |
@@
IV-G Robustness Evaluation
In real-world scenarios, images are inevitably affected by unknown disturbances during transmission and interaction, which pose additional challenges for AI-generated image detection. To investigate robustness under such conditions, we further evaluate various detection methods against common perturbations, including JPEG compression (Quality Factor (QF) = 95, 85, 75) and Gaussian blur ( = 0.5, 1.0, 1.5). As shown in Table VIII and Table VIII, experiments on two different datasets reveal that all methods experience performance degradation under perturbations, weakening their ability to reliably distinguish real from fake images. Despite these challenging conditions, our method consistently outperforms competing approaches, maintaining relatively high accuracy. Notably, its performance remains almost unaffected under varying levels of Gaussian blur, further demonstrating the strong robustness of our approach.
IV-H Comparative Studies of Module Design
@@
1) The reconstructed sliding window attention mechanism greatly improves the ability of attention to capture forgery features. As shown in Table IX, compared with WAttention, our RSWAttention improves the mean Acc. by 24.2%, which suggests that the RSWAttention greatly enhances the attention mechanism to capture fine-grained forgery traces within the local window while modeling the importance and interdependence between local features.
@@
2) Utilizing only the phase component of the FFT can reduce features not related to forgery compared to using the full FFT. As shown in Table X, if we use not only the phase part of the FFT but also the amplitude part as input to the network, the average accuracy of the model drops by 5.2%, confirming that amplitude introduces noise or irrelevant intensity information that confounds the detector. In contrast, the phase, encoding structural details like edges and shapes, aligns with the artifact-relevant features needed for detection. This finding corroborates prior image processing research, where phase information is prioritized over amplitude for structural analysis, reinforcing our design choice to leverage FFT’s phase as a complementary input branch. In addition, replacing FFT with DCT in the complement branch leads to about a 2% decrease in mAcc., as shown in Table XI, demonstrating the effectiveness of relying on FFT-based features.
@@
3) All subbands of DWT are indispensable. Each frequency domain subband of DWT has its own unique meaning, the low-frequency part retains the overall structure and semantic background of the image, and the high-frequency part captures subtle texture and edge information. Deleting any subband will reduce the comprehensiveness of the information in the window. As shown in Table XII, whether using only one frequency domain subband or omitting some frequency domain subbands, the performance of the model will degrade.
4) The window size for each module is optimal. To further validate the effectiveness of our window design, we conduct additional experiments to examine the impact of different window sizes. As shown in Table XIII, any deviation from our chosen configuration results in a noticeable performance drop, which further substantiates the suitability of our design. Specifically, in the FFT-based phase complement branch, each element is derived from the global phase spectrum and thus encodes long-range dependencies and structural relationships among the overall elements. As shown in the first row of table XIII, reducing the attention window restricts interactions to only very local neighborhoods of phase features, weakening the model’s ability to capture global phase relationships and resulting in degraded performance. In contrast, the DWT-based window tiling branch is designed to capture comprehensive and fine-grained spatial-frequency domain information, enabling more detailed feature extraction. After the window tiling, each window displays a highly organized structure: the representation encodes both the four spatially adjacent feature values from the original image and the corresponding feature information from four distinct frequency subbands. This alignment encourages attention to capture coherent dependencies across both neighboring spatial locations and frequency components. Enlarging the tiling window or changing the attention window disrupts this carefully constructed local structure, mixing unrelated spatial regions and subbands within the same window. Consequently, the representation within each window becomes less coherent, leading to a marked decrease in detection performance, as observed in Table XIII.
IV-I Hyperparameter Ablation Study